OpenStudy (anonymous):

PLEASE HELP What is the vertex of a parabola defined by the equation x = 5y2?

OpenStudy (anonymous):

(0.05, 0.05) (0, 0.05) (0, 0) (0, -0.05) Done

OpenStudy (anonymous):

Think of your equation in this fashion, (x-0) = 5(y-0)^2 ...What are the coordinates that you see. What makes x-0 = 0 and what makes y-0 = 0?

OpenStudy (anonymous):

whats the answer?